Seven identical cutes, side 1, 0 m. are arranged a
Answer #1

Ans:-8]Xcom = m1 x1 + m2x2/m1+m2

Xcom = m *0 + m*3 / 2m = 3m/2m =1.5m

Similarly,

Ycom = m*0+ m*2 / 2m = 1m

Zcom = m*0+m*3 /2m = 1.5m

9]Vcom = m1*V1 + m2*V2/m1+m2

              = 12g*3.5 - 8g*2/20g = 1.3m/s

10]Fnet = m*a

The net Fnet force on the ststem is equalto the product of the total mass of the system and the acceleration of the system